OpenStudy (anonymous):

solve sqrt 2x-1 =x-2 . check for extraneous solutions

OpenStudy (helder_edwin):

u have \[ \large \sqrt{2x-1}=x-2 \] how do u get rid of the square root?

OpenStudy (anonymous):

it is fixed that x is greater or equal to 1/2 for the square root to be meaningful. we first fix what values are we allowed to put it in place of x. once we are done with that, solve the equation as how it should be done. squaring on both sides and need to solve the quadratic equation. we get either 2 values of x or only 1 value or maybe no values of x. but the value of x should be greater than 1/2. now try to proceed and find the value of x.

OpenStudy (bradely):

step by step answers posted here http://www.mathskey.com/question2answer/4173/solve-sqrt-2x-1-x-2

OpenStudy (whpalmer4):

Or learn how to get correct answers right here on OpenStudy. On the other site, @bradely claims that x=1 is a solution. Let's check, as the problem specifically tells us to check for extraneous solutions: \[\sqrt{2(1)-1}=(1-2)\]\[\sqrt{1}=-1\] That is incorrect: \(x=1\) is an extraneous solution.

OpenStudy (bradely):

x=5 is the solution